How many pounds of nuts that costs$1.10 per pound must be mixed with 50 pounds of nut that cost $1.80 per pound to make a mixture that cost $1.30 per pound?



Answer by ptaylor(1894)

Let x=number of pounds of nuts that cost $1.10 per pound
Value of these nuts=$1.10x
Then 50-x=number of pounds of nuts that cost $1.80 per pound
Value of these nuts =$1.80(50-x)
Value of final mixture = $1.30(50)
Now we know that the value of the nuts before they are mixed ($1.10x+$1.80(50-x)) is equal to the value of the nuts after they are mixed($1.30*50). So our equation to solve is:
$1.10x+$1.80(50-x)=$1.30*50 get rid of parens
$1.10x+$90-$1.80x=$65 subtract $90 from both sides
$1.10x+$90-$90-$1.80x=$65-$90 combine like terms
-$0.70x=-$25 divide both sides by -$0.70
x=35.714 lbs of nuts that cost $1.10 per pound
CK
(35.714)($1.10)+($1.80)(50-35.714)=$1.30*50
39.2854+25.714=65
65.002~65

Hope this helps---ptaylor